# The Identification Power of

### v1

@@ 7 Conclusion
This paper examines how combining experimental data with observational data improves the identification of DTE parameters. We show that the identified set under the combined data is fully characterized by the joint distribution of each potential outcome, the self-selection variable, and covariates, with the self-selection variable serving as a key source of the identification gain. For a broad class of DTR parameters—represented by super/sub-modular functions and $\varphi$-indicator functions—we derive analytical expressions for the sharp bounds under the combined data. We further characterize necessary and sufficient conditions under which the combined data strictly shrink the identified set, showing that such shrinkage generally occurs unless selection-on-observables holds in the observational data. We also propose a computational approach based on linear programming to compute the sharp bounds, which can incorporate additional structural assumptions such as mutual stochastic monotonicity of the potential outcomes and the generalized Roy model. An empirical application using DRPT data on negative campaign advertisements in a U.S. presidential election illustrates the practical relevance of combining random-assignment and self-selection data sources, and highlights a novel advantage of DRPT designs in improving the informativeness of distributional treatment effects analyses.
@@ Appendix A Numerical Example
This appendix illustrates the identifying power gained by combining experimental and observational data through a simple numerical example. Suppose that our parameter of interest is $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$, the fraction of positive treatment effect. Let $(Y_{1},Y_{0})$ be supported on $1,1]\times[-1,1]$ with the following piecewise uniform density:
$f_{Y_{1}Y_{0}}^{\ast}(y_{1},y_{0})=\begin{cases}\begin{array}[]{c}0.0,\\ 0.5,\\ 0.5,\\ 0.0\end{array}&\begin{array}[]{c}(y_{1},y_{0})\in[-1,0]\times[-1,0]\\ (y_{1},y_{0})\in(0,1]\times[-1,0]\\ (y_{1},y_{0})\in[-1,0]\times(0,1]\\ (y_{1},y_{0})\in(0,1]\times(0,1]\end{array}\end{cases},$ (17)
where $(Y_{1},Y_{0})$ is uniformly distributed at each subrectangle. The true parameter value under this distribution is $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})=0.5$. We do not consider any covariate $X$.
Under the supposed distribution, the marginal distributions $F_{Y_{1}}^{\ast}$ and $F_{Y_{0}}^{\ast}$ are both uniform over [1,1][-1,1]. Proposition 4.2 shows that, under this setting, the sharp identified set for $\theta_{o}$ under experimental data can be computed as $F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$ with $\varphi(y_{1},y_{0})=y_{0}-y_{1}$ and $\delta=0$. It results in $\Theta_{I}=[0,1]$. Thus, experimental data alone is completely uninformative in this example.
We further suppose that $\mathbb{P}(S=1)=1/2$ and that the conditional densities of $(Y_{1},Y_{0})$ given $s=0,1$ have the following piecewise uniform density functions:
@@
with $(a,b)=(0.8,0.2)$. These conditional densities imply that individuals who self-select treatment and no-treatment are more likely to have positive and negative treatment effects, respectively. Note that the assumed conditional densities $f_{Y_{1}Y_{0}|S}$ and self-selection probability $\mathbb{P}(S=\cdot)$ are consistent with the unconditional density (17) (i.e., $f_{Y_{1}Y_{0}}^{\ast}(y_{1},y_{0})=\mathbb{P}(S=1)\cdot f_{Y_{1}Y_{0}|S}^{\ast}(y_{1},y_{0}|1)+\mathbb{P}(S=0)\cdot f_{Y_{1}Y_{0}|0}^{\ast}(y_{1},y_{0}|0)$).
Under the supposed distribution, the conditional CDFs $F_{Y_{d}|S}^{\ast}(y_{1}|s)$ for $d=0,1$ and $s=0,1$ are given by
@@
Then, together with $\mathbb{P}(S=1)=0.5$, the identified set $\Theta_{IC}$ under the combined data can be computed with Proposition 4.2 (ii), and results in $\Theta_{IC}=[0.3,0.7]$. This identified set is substantially tighter than $\Theta_{I}=[0,1]$, indicating that the combination of experimental and observational data considerably improves the identification of $\theta_{o}$.
Now consider another case where $a=1$ and $b=0$ in the conditional densities above. This corresponds to the case that the self-selection $S$ follows the Roy selection model $S=\mathbf{1}\{Y_{1}-Y_{0}>0\}$. The supposed conditional densities $f_{Y_{1}Y_{0}|S}$ and self-selection probability $\mathbb{P}(S=\cdot)$ remain consistent with the unconditional density (17). In this case, the identified set $\Theta_{IC}$, computed by Proposition 4.2 (ii), collapses to the singleton $\Theta_{IC}=\{0.5\}$; hence, point identification is achieved with the combined data while experimental data alone is uninformative (i.e., $\Theta_{I}=[0,1]$). This is notable result because the characterization of $\Theta_{IC}$ does not rely on prior knowledge of the Roy selection mechanism but can achieve point identification through combining experimental and observational data.
@@ Appendix B Linear Programming
This appendix shows that when all random variables are discrete, the optimization problems (13)–(16) can be formulated as a finite-dimensional linear program. For a generic distribution function $F_{Y_{1}Y_{0}SX}$, let $f_{Y_{1}Y_{0}SX}$ denote its density function. Let $f_{Y_{d}SX}^{\ast}$ denote the true density of $(Y_{d},S,X)$ generated by $F_{Y_{d}SX}^{\ast}$, which can be identified as in Lemma 3.1.
The optimization problems (13)–(16) can then be reformulated in terms of density functions as follows:
@@
$\frac{\sum_{y_{d}\leq t,Y_{d^{\prime}}=y}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)}{f_{Y_{d^{\prime}}SX}^{\ast}(y,s,x)}-\frac{\sum_{y_{d}\leq t,y_{d^{\prime}}=y^{\prime}}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)}{f_{Y_{d^{\prime}}SX}^{\ast}(y^{\prime},s,x)}\geq 0$ (21)
@@
$\frac{\sum_{y_{1}-y_{0}>c}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},1,x)}{f_{SX}^{\ast}(S=1,X=x)}-\frac{\sum_{y_{1}-y_{0}>c}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},0,x)}{f_{SX}^{\ast}(S=0,X=x)}\geq 0\ \ \forall t,x$ (22)
Constraint (19) ensures that $f_{Y_{1}Y_{0}SX}$ is a valid probability density function. Constraints (20)–(22) correspond to the original constraints (14)–(16), respectively. Note that that the denominators in constraints (21) and (22) use the true densities $f_{Y_{d^{\prime}}SX}^{\ast}$ and $f_{SX}^{\ast}$, rather than $f_{Y_{d^{\prime}}SX}$ and $f_{SX}$, which is valid because constraint (20) implies $f_{Y_{d}SX}=f_{Y_{d}SX}^{\ast}$ for $d=0,1$.
Since $f_{Y_{1}Y_{0}SX}$ takes finitely many values and all constraints (19)–(22) are linear in $f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)$, the optimization problem (18)–(22) is a finite-dimensional linear program.
Appendix C Proofs
@@
We begin by showing that $\mathbb{P}(S=s\mid X=x)=\mathbb{P}(D=s\mid G=\mathrm{obs},X=x)$. From Assumption 2.2, we have $\mathbb{P}(D=s\mid G=\mathrm{obs},X=x)=\mathbb{P}(S=s\mid G=\mathrm{obs},X=x)$. Then, by Assumption 2.4, it follows that $\mathbb{P}(S=s\mid G=\mathrm{obs},X=x)=\mathbb{P}(S=s\mid X=x)$, establishing the desired equality.
We next consider the identification of $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x)$. When $d=s$, we have
@@
where the second equality follows from Assumption 2.2, and the third from Assumption 2.4. Hence the result (4) for $d=s$ holds.
For the case $d\neq s$, we begin with the following decomposition:
$F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)$ $=\mathbb{P}(S=d\mid X=x)\cdot F_{Y_{d}\mid SX}(\cdot\mid d,x)+\mathbb{P}(S=s\mid X=x)\cdot F_{Y_{d}\mid SX}(\cdot\mid s,x).$
@@
$F_{Y_{d}\mid SX}(\cdot\mid s,x)=\frac{F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)-\mathbb{P}(S=d\mid X=x)\cdot F_{Y_{d}\mid SX}(\cdot\mid d,x)}{\mathbb{P}(S=s\mid X=x)}.$
In this expression, $F_{Y_{d}|X}^{\ast}(\cdot|x)$ is identified as $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$ under Assumptions 2.3 and 2.4, and $\mathbb{P}(S=\cdot|X=x)$ and $F_{Y_{d}|SX}(\cdot|d,x)$ are identified as $\mathbb{P}(S=\cdot|X=x)=\mathbb{P}(D=\cdot|G=\mathrm{obs},X=x)$ and $F_{Y_{d}|SX}^{\ast}(\cdot|d,x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$, respectively, as we have shown above. Therefore, the result (4) for $d\neq s$ holds. ∎
@@
where the first equality follows from Assumption 2.2, and the second from Assumption 2.4. The result (ii) follows the same argument as in the first part of the proof of Lemma 3.1. ∎
@@
We begin by showing that $\Theta_{I}\subseteq\widetilde{\Theta}_{I}$. Fix any $\theta\in\Theta_{I}$. By definition, there exists a distribution function $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ such that $\theta=\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$. Because $F$ satisfies Assumptions 2.12.5 (with $F^{\ast}$ replaced by $F$), and because $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ and $F_{X}=F_{X}^{\ast}$, we have that for $d=0,1$ and all $(y,x)$,
$F_{Y_{d}X}(y,x)$ $=F_{Y_{d}|X}(y|x)\cdot F_{X}(x)=F_{Y|DGX}(y|d,\mathrm{exp},x)\cdot F_{X}(x)$
$=F_{Y|DGX}^{\ast}(y|d,\mathrm{exp},x)\cdot F_{X}^{\ast}(x)=F_{Y_{d}|X}^{\ast}(y|x)\cdot F_{X}^{\ast}(x)$
$=F_{Y_{d}X}^{\ast}(y,x),$ (23)
where the second and fourth equalities follow from Assumptions 2.1, 2.3, and 2.4, and the third follows from the equations $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$ and $F=F^{\ast}$.
Using Sklar’s Theorem, we know that there exists a conditional copula function $C(\cdot,\cdot|x)\in\mathcal{C}$ such that, for almost all $x$, the conditional joint distribution $F_{Y_{1}Y_{0}|X}$ can be expressed as
@@
where the fifth line follows from (23), and the last line follows from the definition of $\widetilde{\Theta}_{I}$. We thus have $\theta\in\widetilde{\Theta}_{I}$. Since this result holds for any $\theta\in\Theta_{I}$, we have $\Theta_{I}\subseteq\widetilde{\Theta}_{I}$.
We next show that $\widetilde{\Theta}_{I}\subseteq\Theta_{I}$. Fix any $\theta\in\widetilde{\Theta}_{I}$. By definition of $\widetilde{\Theta}_{I}$, there exists a conditional copula function $C(\cdot,\cdot|x)\in\mathcal{C}$ such that
@@
We will show that there exists a distribution function $F$ in $\mathcal{F}_{\mathrm{exp}}^{\ast}$ that reproduces $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ in the following way:
$F=F_{Y|Y_{1}Y_{0}DSGX}\times F_{S|Y_{1}Y_{0}DGX}\times F_{Y_{1}Y_{0}|DGX}\times F_{DGX},$
where
@@
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{exp},X)=F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{obs},X);$ (28)
$F_{Y|Y_{1}Y_{0}DSGX}(y)=\mathbf{1}\{DY_{1}+(1-D)Y_{0}\leq y\}$ (29)

### v4

@@ 7 Conclusion
This study investigates how combining experimental and observational data can improve the identification of DTE parameters. We show that the identified set under the combined data is fully characterized by the joint distribution of each potential outcome, latent self-selection variable, and covariates, with latent self-selection serving as the key source of additional identification power. For a broad class of DTE parameters represented by super/sub-modular functions and $\varphi$-indicator functions, we derive sharp bounds under the combined data. We further establish necessary and sufficient conditions under which the data combination strictly shrinks the identified set, suggesting that such shrinkage generally occurs unless selection-on-observables holds in the observational data. We also propose a linear programming approach for computing sharp bounds while incorporating additional structural assumptions, such as positive dependence of potential outcomes and generalized Roy model selection. An empirical application using DRPT data on negative campaign advertisements in the U.S. presidential election illustrates the practical value of combining random-assignment and self-selection data, and highlights a novel advantage of DRPT designs.
Table
Table 1: Estimated Identified Sets of $\mathbb{P}(Y_{1}<Y_{0})$ for John McCain and Barack Obama
Panel (a): John McCain
Inclusion of Population
Self-selection Asm.PD All Democrats Republicans Selected T Selected NT
No No $[0.13,0.90]$ $[0.12,0.88]$ $[0.11,0.92]$
Yes No $[0.21,0.81]$ $[0.20,0.81]$ $[0.20,0.84]$ $[0.36,0.88]$ $[0.11,0.76]$
No Yes $[0.17,0.79]$ $[0.16,0.79]$ $[0.15,0.80]$
Yes Yes $[0.40,0.53]$ $[0.33,0.56]$ $[0.45,0.60]$ $[0.46,0.56]$ $[0.36,0.52]$
Panel (b): Barack Obama
Inclusion of Population
Self-selection Asm.PD All Democrats Republicans Selected T Selected NT
No No $[0.17,0.89]$ $[0.12,0.88]$ $[0.17,0.87]$
Yes No $[0.22,0.77]$ $[0.13,0.73]$ $[0.27,0.81]$ $[0.30,0.65]$ $[0.18,0.84]$
No Yes $[0.18,0.72]$ $[0.12,0.74]$ $[0.17,0.70]$
Yes Yes $[0.38,0.56]$ $[0.43,0.57]$ $[0.26,0.56]$ $[0.35,0.43]$ $[0.41,0.63]$
  • Notes: Each panel reports the estimated identified sets of $\mathbb{P}(Y_{1}<Y_{0})$ for each candidate, covering the full population and specified subpopulations (Democrats, Republicans, and each self-selection type $s\in\{0,1\}$) across four combinations: (i) including vs. excluding the self-selection sample ($G=\mathrm{obs}$) and (ii) imposing vs. not imposing Assumption 5.1 (Positive Dependence). “Self-selection” indicates whether the self-selection sample is included, and “Asm. PD” indicates whether Assumption 5.1 is imposed. “Democrats” and “Republicans” refer to self-identified party affiliation. “Selected T” and “Selected NT” denote individuals who self-selected into viewing and not viewing the negative campaign advertisements, respectively.
@@ Appendix A Numerical Example
This appendix illustrates the identifying power gained by combining experimental and observational data through a simple numerical example. Suppose that the parameter of interest is $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})$. Let $(Y_{1},Y_{0})$ be supported on $1,1]\times[-1,1]$ with the following piecewise uniform joint density:
$f_{Y_{1}Y_{0}}^{\ast}(y_{1},y_{0})=\begin{cases}\begin{array}[]{c}0.0,\\ 0.5,\\ 0.5,\\ 0.0,\end{array}&\begin{array}[]{c}(y_{1},y_{0})\in[-1,0]\times[-1,0]\\ (y_{1},y_{0})\in(0,1]\times[-1,0]\\ (y_{1},y_{0})\in[-1,0]\times(0,1]\\ (y_{1},y_{0})\in(0,1]\times(0,1]\end{array}\end{cases},$ (19)
where $(Y_{1},Y_{0})$ is uniformly distributed within each subrectangle. Under this distribution, the true parameter value is $\theta_{o}=\mathbb{P}(Y_{1}>Y_{0})=0.5$. We do not introduce any covariate $X$.
Under the supposed distribution, the marginal distributions $F_{Y_{1}}^{\ast}$ and $F_{Y_{0}}^{\ast}$ are both uniform on [1,1][-1,1]. By Proposition 4.3, the sharp identified set for $\theta_{o}$ under the experimental data can be computed as $F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$ with $\varphi(y_{1},y_{0})=y_{0}-y_{1}$ and $\delta=0$, yielding $\Theta_{I}=[0,1]$. Hence, in this example, the experimental data alone provide no information about $\theta_{o}$.
We further suppose that $\mathbb{P}(S=1)=1/2$ and that the conditional joint densities of $(Y_{1},Y_{0})$ given $S=0,1$ are piecewise uniform as follows:
@@
with $(a,b)=(0.8,0.2)$. These conditional densities imply that individuals who self-select into treatment are more likely to experience positive treatment effects, whereas those who self-select into no-treatment are more likely to experience negative treatment effects. Note that the assumed conditional densities $f_{Y_{1}Y_{0}|S}$ and the self-selection probability $\mathbb{P}(S=\cdot)$ are consistent with the unconditional density in (19) (i.e., $f_{Y_{1}Y_{0}}^{\ast}(y_{1},y_{0})=\sum_{s=0,1}\mathbb{P}(S=s)\cdot f_{Y_{1}Y_{0}|S}^{\ast}(y_{1},y_{0}|s)$).
Under the assumed distribution, the conditional CDFs $F_{Y_{d}|S}^{\ast}(y_{1}|s)$ for $d=0,1$ and $s=0,1$ are given by
@@
Together with $\mathbb{P}(S=1)=0.5$, the identified set $\Theta_{IC}$ under the combined data can be computed using Proposition 4.3(ii), yielding $\Theta_{IC}=[0.3,0.7]$. This interval is substantially narrower than $\Theta_{I}=[0,1]$, demonstrating that combining experimental and observational data can substantially improve the identification of $\theta_{o}$.
Now consider the case in which $a=1$ and $b=0$ in the conditional densities above. This corresponds to a Roy selection model, in which self-selection is determined by $S=\mathbf{1}\{Y_{1}-Y_{0}>0\}$. Under this specification, the conditional densities $f_{Y_{1}Y_{0}|S}^{\ast}$ and the self-selection probability $\mathbb{P}(S=\cdot)$ remain consistent with the unconditional density in (19). In this case, the identified set $\Theta_{IC}$, computed using Proposition 4.3(ii), collapses to the singleton $\Theta_{IC}=\{0.5\}$. Thus, the combined data achieve point identification, whereas the experimental data alone remain completely uninformative ($\Theta_{I}=[0,1]$). Notably, this point identification is attained without any prior knowledge of the Roy selection mechanism; rather, it arises solely from combining experimental and observational data.
@@ Appendix B Linear Programming
This appendix shows that when all random variables are discrete, the optimization problems (15)–(18) can be formulated as a finite-dimensional linear program. For a generic distribution function $F_{Y_{1}Y_{0}SX}$, let $f_{Y_{1}Y_{0}SX}$ denote the density function. Let $f_{Y_{d}SX}^{\ast}$ denote the true density of $(Y_{d},S,X)$ induced by $F_{Y_{d}SX}^{\ast}$, which can be identified as in Lemma 3.2.
The optimization problems (15)–(18) can then be reformulated in terms of the density functions as follows:
@@
$\frac{\sum_{y_{d}\leq t,Y_{d^{\prime}}\leq y}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)}{\sum_{y_{d^{\prime}}\leq y}f_{Y_{d^{\prime}}SX}^{\ast}(y_{d^{\prime}},s,x)}-\frac{\sum_{y_{d}\leq t,y_{d^{\prime}}\leq y^{\prime}}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)}{\sum_{y_{d^{\prime}}\leq y^{\prime}}f_{Y_{d^{\prime}}SX}^{\ast}(y_{d^{\prime}},s,x)}\geq 0$ (23)
@@
$\frac{\sum_{y_{1}-y_{0}>c}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},1,x)}{f_{SX}^{\ast}(1,x)}-\frac{\sum_{y_{1}-y_{0}>c}f_{Y_{1}Y_{0}SX}(y_{1},y_{0},0,x)}{f_{SX}^{\ast}(0,x)}\geq 0\ \ \forall c,x.$ (24)
The constraint in equation (21) ensures that $f_{Y_{1}Y_{0}SX}$ is a valid probability density function, while equations (22)–(24) correspond to the original constraints in (16)–(18), respectively. Note that the denominators in (23) and (24) involve the true densities $f_{Y_{d^{\prime}}SX}^{\ast}$ and $f_{SX}^{\ast}$, rather than $f_{Y_{d^{\prime}}SX}$ and $f_{SX}$. This substitution is valid because (22) implies $f_{Y_{d}SX}=f_{Y_{d}SX}^{\ast}$ for $d\in\{0,1\}$.
Since $f_{Y_{1}Y_{0}SX}$ takes finitely many values and all constraints in (21)–(24) are linear in $f_{Y_{1}Y_{0}SX}(y_{1},y_{0},s,x)$, the optimization problem (20)–(24) is a finite-dimensional linear program.
The following remark explains how we estimate the identified sets under Assumption 5.1 in the empirical application presented in Section 6.
Remark B.1 (Empirical Application in Section 6).
In our empirical application in Section 6, when Assumption 5.1 is imposed together with the self-selection sample, we estimate the sharp bounds by solving the linear program (20)–(23), setting $\psi(y_{1},y_{0})=\mathbf{1}\{y_{1}<y_{0}\}$ and replacing $F_{Y_{d}SX}^{\ast}$ and $f_{SX}^{\ast}$ with their empirical counterparts constructed according to Lemma 3.2.151515Specifically, each component in equations (4) and (5) is estimated using the corresponding empirical distribution or density. [^16] For equation (23), which is derived from Assumption 5.1, we exclude $t$- and $y$-values below the 2.5th percentile and above the 97.5th percentile of the estimated outcome distribution to mitigate boundary bias. To reduce memory usage and computation time, we additionally thin the grid by retaining every third $t$- and $y$-value. We do not impose the constraint in equation (24) because Assumption 5.2 is not maintained.
When Assumption 5.1 is imposed without the self-selection sample, we estimate the sharp bounds analogously using the corresponding linear program, but excluding the self-selection variable. Specifically, we drop the self-selection variable $s$ from the linear program (20)–(23) and replace $F_{Y_{d}X}^{\ast}$ and $f_{X}^{\ast}$ with their empirical counterparts based on the sample with $G=\mathrm{exp}$. [^17] That is, estimation relies solely on the experimental subsample.
We solve the linear programs using Gurobi. In finite samples, the optimization problem may become infeasible due to conflicting constraints. In such cases, we apply Gurobi’s feasrelax procedure to constraint (23), which introduces slack variables and solves an auxiliary optimization problem that minimizes the total relaxation required. This approach restores feasibility while keeping deviations from the original constraints as small as possible.
When estimating $\mathbb{P}(Y_{1}<Y_{0}\mid X=x)$, we restrict the sample to units with $X=x$. To estimate $\mathbb{P}(Y_{1}<Y_{0}\mid S=s)$, following the discussion in Section 2.2, we define $\psi(y_{1},y_{0})=\mathbf{1}\{y_{0}>y_{1}\}\cdot\mathbf{1}\{S=s\}/\mathbb{P}(S=s)$, where $\mathbb{P}(S=s)$ is estimated by the sample fraction of $D=s$ among units with $G=\mathrm{obs}$.
Appendix C Proofs and a Preliminary Lemma
@@
We begin by showing equation (4). By Assumption 2.2,
$\mathbb{P}(D=s\mid G=\mathrm{obs},X=x)=\mathbb{P}(S=s\mid G=\mathrm{obs},X=x).$
Then, by Assumption 2.4, it follows that
$\mathbb{P}(S=s\mid G=\mathrm{obs},X=x)=\mathbb{P}(S=s\mid X=x).$
Combining these results yields equation (4).
We next consider the identification of $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x)$, thereby showing equation (5). When $d=s$, we have
@@
where the second equality follows from Assumption 2.2, and the third from Assumption 2.4. Hence, equation (5) hols for $d=s$.
For the case $d\neq s$, we begin with the decomposition:
$F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)$ $=\mathbb{P}(S=d\mid X=x)\cdot F_{Y_{d}\mid SX}^{\ast}(\cdot\mid d,x)+\mathbb{P}(S=s\mid X=x)\cdot F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x).$
@@
$F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x)=\frac{F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)-\mathbb{P}(S=d\mid X=x)\cdot F_{Y_{d}\mid SX}^{\ast}(\cdot\mid d,x)}{\mathbb{P}(S=s\mid X=x)}.$
In this expression, $F_{Y_{d}|X}^{\ast}(\cdot|x)$ is identified as $F_{Y_{d}|X}^{\ast}(\cdot|x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)$ under Assumptions 2.3 and 2.4. Moreover, $\mathbb{P}(S=\cdot|X=x)$ and $F_{Y_{d}|SX}(\cdot|d,x)$ are identified as $\mathbb{P}(S=\cdot|X=x)=\mathbb{P}(D=\cdot|G=\mathrm{obs},X=x)$ and $F_{Y_{d}|SX}^{\ast}(\cdot|d,x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{obs},x)$, as shown above. Therefore, equation (5) holds for $d\neq s$. ∎
@@
where the first equality follows from Assumption 2.2 and the second from Assumption 2.4. The result (ii) follows by the same argument as in the first part of the proof of Lemma 3.2. ∎
@@
We begin by showing that $\Theta_{I}\subseteq\widetilde{\Theta}_{I}$. Fix any $\theta\in\Theta_{I}$. By definition, there exists a distribution function $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ such that $\theta=\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$. Since $F$ satisfies Assumptions 2.12.5 (with $F^{\ast}$ replaced by $F$) and $F_{YDX|G}(\cdot,\cdot,\cdot|\mathrm{exp})=F_{YDX|G}^{\ast}(\cdot,\cdot,\cdot|\mathrm{exp})$, it follows that for each $d\in\{0,1\}$ and all $x$,
$F_{Y_{d}|X}(\cdot|x)=F_{Y|DGX}(\cdot|d,\mathrm{exp},x)=F_{Y|DGX}^{\ast}(\cdot|d,\mathrm{exp},x)=F_{Y_{d}|X}^{\ast}(\cdot|x),$ (25)
where the first and third equalities follow from Assumptions 2.1, 2.3, and 2.4, and the second from $F_{YDX\mid G}(\cdot,\cdot,\cdot\mid\mathrm{exp})=F_{YDX\mid G}^{\ast}(\cdot,\cdot,\cdot\mid\mathrm{exp})$.
By Sklar’s Theorem (e.g., Nelsen, 2006, Theorem 2.3.3), there exists a conditional copula function $C(\cdot,\cdot\mid x)\in\mathcal{C}$ such that, for almost all $x$, the conditional joint distribution $F_{Y_{1}Y_{0}\mid X}$ can be written as
@@
where the fifth line follows from (25) and the condition $F_{X}=F_{X}^{\ast}$ in $\mathcal{F}_{\mathrm{exp}}^{\ast}$, and the last line follows from the definition of $\widetilde{\Theta}_{I}$. Hence, $\theta\in\widetilde{\Theta}_{I}$. Since this argument holds for any $\theta\in\Theta_{I}$, we conclude that $\Theta_{I}\subseteq\widetilde{\Theta}_{I}$.
We next show that $\widetilde{\Theta}_{I}\subseteq\Theta_{I}$. Fix any $\theta\in\widetilde{\Theta}_{I}$. By the definition of $\widetilde{\Theta}_{I}$, there exists a conditional copula function $C(\cdot,\cdot|x)\in\mathcal{C}$ such that
@@
We now show that there exists a distribution function $F\in\mathcal{F}_{\mathrm{exp}}^{\ast}$ that reproduces $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ hierarchically, defined by
@@
$F_{S|Y_{1}Y_{0}DGX}(s|Y_{1},Y_{0},D,\mathrm{exp},X)=F_{S|Y_{1}Y_{0}GX}(s|Y_{1},Y_{0},\mathrm{obs},X);$ (30)
$F_{Y|Y_{1}Y_{0}DSGX}(y)=\mathbf{1}\{DY_{1}+(1-D)Y_{0}\leq y\}.$ (31)